OpenStudy (anonymous):

Find the slope of the line. y = -3x

OpenStudy (anonymous):

it's the number in front of the x

OpenStudy (lexipex61):

The answer is -3.

OpenStudy (anonymous):

oh thanks...but what if there was a number infront of y too

OpenStudy (anonymous):

You can isolate the y

OpenStudy (anonymous):

2y = 3x gives you \[y = \frac{ 3 }{ 2 }x\]

OpenStudy (anonymous):

solve for y and it will be the number in front of x. ex. 2x + 3y = 6 solve for y 3y = -2x + 6 => \(y = \frac{-2}{3}x +2\)

OpenStudy (anonymous):

the slope is \(\frac{-2}{3}\)

OpenStudy (anonymous):

thanks